# Automated Deep Learning Optimization via DSL-Based Source Code Transformation

### v1

@@ 3.5. Synthesis-based Code Transformation
| 1 | BertSelfAttention | BertSelfAttention with xformers’ attention ops | xformers (xsf, 2023) |
| 2 | T5Attention | T5Attention with xformers’ attention ops | xformers (xsf, 2023) |
| 3 | GPT2Attention | GPT2Attention with xformers’ attention ops | xformers (xsf, 2023) |
| 4 | softmax | A drop-in replacement to torch.nn.softmax | triton (noa, 2023b) |
| 5 | Dropout_LayerNorm | Fusing Dropout and LayerNorm | epoi (epo, 2023) |
| 6 | biased_GeLU | Fusing biased Linear and GeLU activation | epoi (epo, 2023) |
| 7 | Conv_BatchNorm | Fusing Conv2d and BatchNorm2d | PyTorch (pyt, 2023b) |
| 8 | Linear_BatchNorm | Fusing Linear and BatchNorm1d | PyTorch (pyt, 2023b) |
| 9 | fused_QKV | Fusing three Linear layers as q, k, and v in encoder | slapo (noa, 2023a) |

### v2

@@ 3.5. Synthesis-based Code Transformation
1 | BertSelfAttention | BertSelfAttention with xformers’ attention ops | xformers (xsf, 2023) |
2 | T5Attention | T5Attention with xformers’ attention ops | xformers (xsf, 2023) |
3 | GPT2Attention | GPT2Attention with xformers’ attention ops | xformers (xsf, 2023) |
4 | softmax | A drop-in replacement to torch.nn.softmax | triton (noa, 2023b) |
5 | Dropout_LayerNorm | Fusing Dropout and LayerNorm | epoi (epo, 2023) |
6 | biased_GeLU | Fusing biased Linear and GeLU activation | epoi (epo, 2023) |
7 | Conv_BatchNorm | Fusing Conv2d and BatchNorm2d | PyTorch (pyt, 2023b) |
8 | Linear_BatchNorm | Fusing Linear and BatchNorm1d | PyTorch (pyt, 2023b) |
9 | fused_QKV | Fusing three Linear layers as q, k, and v in encoder | slapo (noa, 2023a) |